OpenStudy (anonymous):

HELPP FAST!!! PLEASSEEEE find the polynomial that has root -3 then 4 with of multiplicity of 2? ill give medals to anyone who helps me

OpenStudy (anonymous):

start with ( x + 3) ( x - 4)^2 and then multiply out i would cheat

OpenStudy (anonymous):

FOIL? right

OpenStudy (anonymous):

foil schmoil i would wolf http://www.wolframalpha.com/input/?i=%28x%2B3%29%28x-4%29^2

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

i need to find a polynomial

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

look at the link i sent it multiplied out for you

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

our you can do it by hand (x+3)(x-4)(x-4)

OpenStudy (anonymous):

did you click on the link? it has the answer in it

OpenStudy (anonymous):

( x + 3)(x -4)(x-4) = (x + 3)(x^2 - 8x +16) = etc

OpenStudy (anonymous):

thank you